# Twin branching in shape memory alloys: a 1D model with energy dissipation effects

### v3

@@ 5 Discussion
Following Seiner et al. (2020), the specific model discussed above corresponds to the situation in which the twinned domain is bounded by the A–MM interface on its one end and by a free surface on the other end. This is a special case that may be encountered in a single crystal undergoing martensitic transformation. A different situation is expected in the case of a twinned martensite plate formed within a grain in a polycrystalline material. One possible scenario would be that the twinned martensite plate is bounded by two A–MM interfaces (no free surfaces). The proposed approach is directly applicable in such a case as well. It suffices to consider the domain $x\in[-L,L]$ with an adequate boundary condition at $x=-L$, namely $h_{,x}\big{|}_{x=-L}=\bar{\Gamma}/A$, cf. Eq. (16)2. Indeed, the corresponding solution can be obtained directly from the solution obtained for the original problem formulated in Section 3.2 by simply exploiting the symmetry with respect to $x=0$ (note the homogeneous natural boundary condition at $x=0$, Eq. (16)1).
@@ Appendix B Estimation of local interface propagation speed
Figure 1: Estimating the local interface propagation speed $\hat{v}_{\rm n}$: (a) $\alpha=1$; (b) $\alpha\neq 1$. The location of the cell at an earlier instant is indicated in gray.
@@ Appendix C Computational treatment of the evolution problem
The minimization problem (31) is non-smooth, and its computational treatment relies on the augmented Lagrangian technique Alart and Curnier (1991), borrowed here from computational contact mechanics. An equivalent smooth saddle-point problem is thus formulated,
@@
where $\lambda$ is the Lagrange multiplier field that is introduced through the functional ${\cal L}$,
${\cal L}=\int_{0}^{L}\frac{r}{h}\,l(\dot{h},\lambda){\rm{d}}x.$ (54)
@@
The Lagrange functional $L$ is smooth (because function $l(\dot{h},\lambda)$ is continuously differentiable) hence the solution of the saddle-point problem (53) can be found by solving the condition of stationarity of $L$ with respect to arbitrary variations $\delta\dot{h}$ and $\delta\lambda$, which reads

### v4

@@ 5 Discussion
The model developed above is based on a simple estimate of the interfacial energy contribution, Eq. (8), which leads to a particularly simple form of the corresponding term in the free energy density (11). This form corresponds to the assumption that the deviation of the orientation of the actual twin interfaces from the theoretical (nominal) orientation is small and thus negligible and, secondly, the local interfacial energy density is not affected by such small changes of the interface orientation. It is of interest to examine the validity of these assumptions.
Referring to Fig. 1, the change of the interface orientation due to branching is characterized by angle $\beta$ which depends on the spatial derivative $h_{,x}$ of the twin spacing $h$, see Eq. (52). The magnitude of the derivative $h_{,x}$ is the largest at $x=L$, see Fig. 2, and is then equal to $\bar{\Gamma}=\Gamma/A$, according to the boundary condition (17)2. For the specific parameters adopted in Section 3.4, we have $\beta=5^{\circ}$ for $\bar{\Gamma}=1.2$ and $\beta=1.5^{\circ}$ for $\bar{\Gamma}=0.35$ so that the deviation of the orientation is indeed small. The corresponding change of the interface area is even smaller because the length of two twin interfaces (the upper ones in Fig. 1(b)) increases, while the length of the other two interfaces decreases so that the net change is a second-order effect for small $\beta$.
Quantitative assessment of the effect of interface orientation on the interfacial energy is currently not possible since respective data is missing, and actually only rough estimates of the twin boundary energy itself are used. Here, atomistic simulations hold promise to deliver more reliable data, see e.g. recent results for NiTi (La Rosa and Maresca, 2024). On the other hand, in a purely phenomenological manner, the related effects can be estimated using Eq. (7). Adopting Eq. (7) instead of the simplified form in Eq. (8) and using Eqs. (4)2 and (9), the free energy density can be derived in the following extended form,
$\phi(h,h_{,x})=\frac{1}{2}A(h_{,x})^{2}+\frac{2\gamma}{\alpha h}\left(1-\frac{3}{4}(1-\lambda)\lambda\alpha h_{,x}+\frac{9}{16}(1-\lambda)^{2}(1+\lambda)\alpha^{2}(h_{,x})^{2}\right),$ (40)
as a generalization of Eq. (11). The additional terms increase the interfacial energy contribution at $x=L$ by 14% for $\bar{\Gamma}=1.2$ and by 4% for $\bar{\Gamma}=0.35$, while the increase is smaller in the remaining part of the domain (the magnitude of $h_{,x}$ is the largest at $x=L$). However, the overall effect on the twin spacing is insignificant with the difference of less than 1%, except in the vicinity of $x=L$, as our additional computations indicate (the finite-element implementation of the above extended model is straightforward and is not discussed here). Concluding, the effect of interface orientation on the interfacial energy can be disregarded.
Following Seiner et al. (2020), the specific model discussed in Sections 3 and 4 corresponds to the situation in which the twinned domain is bounded by the A–MM interface on its one end and by a parallel free surface on the other end (a non-parallel free surface can also be considered, as discussed below). This is a special case that may be encountered in a single crystal undergoing martensitic transformation. A different situation is expected in the case of a twinned martensite plate formed within a grain in a polycrystalline material. One possible scenario would be that the twinned martensite plate is bounded by two A–MM interfaces (no free surfaces). The proposed approach is directly applicable in such a case as well. It suffices to consider the domain $x\in[-L,L]$ with an adequate boundary condition at $x=-L$, namely $h_{,x}\big{|}_{x=-L}=\bar{\Gamma}/A$, cf. Eq. (17)2. Indeed, the corresponding solution can be obtained directly from the solution obtained for the original problem formulated in Section 3.2 by simply exploiting the symmetry with respect to $x=0$ (note the homogeneous natural boundary condition at $x=0$, Eq. (17)1).
@@
The assumption that the bounding surfaces are parallel, so that the length $L$ is independent of the position within the A–MM interface, can be relaxed, at least when the deviation is not too big. In such a case, the twin spacing $h$ would depend not only on the distance from the A–MM interface but also on the position within the A–MM interface, including in the third (out-of-plane) direction. Recall that the model assumes that there is no periodicity nor perfect arrangement of branching generations, hence individual twin lamellae branch independently and $h$ is an average twin spacing. The energetic cost of the additional in-plane variation of $h$ is thus expected to be insignificant. Accordingly, the 1D model can be solved locally at each position within the A–MM interface using the corresponding local length $L$. In practical terms, care would be needed here, since such a local 1D model would describe branching along an axis inclined with respect to the A–MM interface (parallel to the twin interfaces), while in the present formulation $L$ is measured in the direction normal to the A–MM interface. This is, however, only a technical detail.
@@ Appendix B Estimation of local interface propagation speed
Figure 1: Estimating the local interface propagation speed $\hat{v}_{\rm n}$: (a) $\alpha=1$; (b) $\alpha\neq 1$. The location of the cell at an earlier instant is indicated by gray lines.
@@ Appendix C Computational treatment of the evolution problem
The minimization problem (32) is non-smooth, and its computational treatment relies on the augmented Lagrangian technique (Alart and Curnier, 1991), borrowed here from computational contact mechanics. An equivalent smooth saddle-point problem is thus formulated,
@@
where $\lambda$ denotes (in this appendix only) the Lagrange multiplier field that is introduced through the functional $\Lambda$,
$\Lambda=\int_{0}^{L}\frac{r}{h}\,l(\dot{h},\lambda){\rm{d}}x.$ (56)
@@
The Lagrange functional ${\cal L}$ is smooth (because function $l(\dot{h},\lambda)$ is continuously differentiable) hence the solution of the saddle-point problem (55) can be found by solving the condition of stationarity of ${\cal L}$ with respect to arbitrary variations $\delta\dot{h}$ and $\delta\lambda$, which reads